# Characterisation of the big Ramsey degrees of the generic partial order

### v3

@@ 7. Upper bounds
Given $S\subseteq\Sigma^{*}$ we call function $f\colon S\to\Sigma^{*}$ shape-preserving if $\tau_{S}(w)=\tau_{f[S]}(f(w))$ for every $w\in S$.
@@
    For every $g\in\mathrm{Shape}\allowbreak(S,\allowbreak\Sigma^{*})$ there exists $w\in\Pi^{*}_{\max(S)}$ such that for every $0\leq i<n$ it holds that $g(u^{i})=g(u^{i}|_{k-1})^{\frown}{u^{i}_{k-1}}^{\frown}{w^{0}_{i}}^{\frown}\cdots^{\frown}w^{|w|-1}_{i}$.
@@
$I=\begin{cases}0\quad&\text{if }i=K+1,\\ \min\{I<\omega:W_{I}=\lambda_{i-K-2}\}\quad&\text{if }i>K+1.\end{cases}$
@@ Acknowledgement
We would like to thank to the anonymous referee for remarks which improved clarity of this paper and to Štěpán Vodseďálek for pointing out important typo in Definition 1.5.

### v4

@@ 7. Upper bounds
Given $S\subseteq\Sigma^{*}$ we call function $f\colon S\to\Sigma^{*}$ shape-preserving if $\tau_{S}(w)=\tau_{f[S]}(f(w))$ (see Definition 6.4) for every $w\in S$.
@@
    For every $g\in\mathrm{Shape}\allowbreak(S,\allowbreak\Sigma^{*})$ there exists $w\in\Pi^{*}_{\max(S)}$ such that for every $0\leq i<n$ it holds that $g(u^{i})=g^{0}(u^{i}|_{k-1})^{\frown}{u^{i}_{k-1}}^{\frown}{w^{0}_{i}}^{\frown}\cdots^{\frown}w^{|w|-1}_{i}$.
@@
$I=\begin{cases}0\quad&\text{if }i=K+1,\\ \min\{I^{\prime}<\omega:W_{I^{\prime}}=\lambda_{i-K-2}\}\quad&\text{if }i>K+1.\end{cases}$
@@ 9.1. Comparison to big Ramsey degrees of the order of rationals
Remark 9.2.
It is interesting to note the differences between Definition 1.5 and Definition 9.1. Since in linear orders there are no incomparable pairs, we only need a two-letter alphabet (there is no need for $X$). Moreover, there are only two interesting events: splitting and leaf. Similarly to Remark 3.7, levels of Devlin types can again be seen as approximations of the structure (this time a linear order) represented by the Devlin type. However, these approximations are quite easy: they are again linear orders given by the lexicographic ordering where a vertex of the approximation corresponds to a linear interval of the final order. We thus do not obtain a new kind of level-structures.
@@ 9.2. The triangle-free graph
Remark 9.7.
Tringle-free diares use 00 to represent non-edges and $1$ to represent edges using the “passing number” representation which has been known for the Rado graph [41, 27]. It is again interesting to understand triangle-free diaries in the context of their level structures. While for the Rado graph the level structure is trivial and holds no information except for the lexicographic order, in the triangle-free graph it distinguishes two types of vertices. If vertex has no neighbour defined yet (it is a word $0\cdots 0$), it may split into vertices connected by an edge. There is at most one such vertex in a level strcture. On the other hand, if vertex has a neighbour already defined, splitting it to two vertices connected by an edge would close a triangle. Consequently, it is an interesting event when the first neighbour of a vertex is introduced. Moreover, a given pair of vertices may or may not have a common neighbour. If common neighbour exists, there will be no edges between their respective successors. This is recorded by relation $\perp$ and the corresponding interesting event.
@@
10. Concluding remarks
The techniques presented in this paper are quite general and can be applied to a relatively wide class of structures in finite relational languages with relations of arity at most 2, defined by special triangle conditions isolated in [5]. Important examples of such structures are metric spaces with finitely many distances, or metric spaces associated to primitive metrically homogeneous graphs of finite diameter from Cherlin’s catalogue of metrically homogeneous graphs [12, 1].
However, forbidding a configuration on 4 or more vertices (such as in the universal homogeneous $K_{4}$-free graph) will run into a limitation of regular trees used here. This can be overcome in two ways. Either a Ramsey theorem for coding trees can be proved [17, 19, 44, 3] or a theorem for trees with finite but unbounded branching can be used [4] to represent a more regular tree constructed from a kind of amalgamation of all coding trees corresponding to all possible enumerations of the structure, as used in this paper.
Structures with relations of arity 3 and more require trees with unbounded branching together with more involved trees of weak types. Upper bounds for structures with very simple restrictions can be obtained using the product form of the Milliken tree theorem [7, 10], while more general results can be obtained using trees with a successor operation [4], see also the announcement in [2]. Even the characterisation of big Ramsey degrees of 3-uniform hypergraphs shows new and somewhat surprising phenomena, and is a topic of a follow-up paper.
@@ Acknowledgement
We would like to thank the two anonymous referees for remarks which significantly improved clarity of this paper, and to Štěpán Vodseďálek for pointing out an important typo in Definition 1.5.